# Learning Successor Features with Distributed Hebbian Temporal Memory

### v3

@@ Abstract
This paper presents a novel approach to address the challenge of online temporal memory learning for decision-making under uncertainty in non-stationary, partially observable environments. The proposed algorithm, Distributed Hebbian Temporal Memory (DHTM), is based on factor graph formalism and a multicomponent neuron model. DHTM aims to capture sequential data relationships and make cumulative predictions about future observations, forming Successor Features (SF). Inspired by neurophysiological models of the neocortex, the algorithm utilizes distributed representations, sparse transition matrices, and local Hebbian-like learning rules to overcome the instability and slow learning process of traditional temporal memory algorithms like RNN and HMM. Experimental results demonstrate that DHTM outperforms LSTM and a biologically inspired HMM-like algorithm, CSCG, in the case of non-stationary datasets. Our findings suggest that DHTM is a promising approach for addressing the challenges of online sequence learning and planning in dynamic environments.
@@ 1 Introduction
Modelling sequential data is one of the essential tasks in artificial intelligence as it has many applications, including decision-making and learning world models [1], natural language processing [2], conversational AI [3], time series analysis [4], and video and music generation [5]. One of the classical approaches to modelling sequential data is to form a representation that stores and condenses the most relevant information about a sequence, and to find a general transformation rule of this information through the dimension of time [6, 7, 8]. We refer to the class of algorithms that use this approach as Temporal Memory (TM) algorithms because they essentially model the cognitive ability of complex living organisms to remember the past experience and make future predictions based on that memory [9, 10, 11, 12].
@@
This paper addresses the problem of online sequence learning for planning and decision-making under uncertainty, which can be formalized as reinforcement learning (RL) for a partially observable Markov decision process (POMDP) [13, 14]. Inferring the hidden state in a partially observable environment is a sequence modelling problem, since it requires processing a sequence of observations to obtain enough information about the hidden state. One of the most efficient representations of hidden states for discrete POMDPs is the Successor Representation (SR), which disentangles hidden states and goals given by the reward function [15, 16, 17]. The Successor Features framework is an extension of SR to continuous POMDP, using the same idea of value function decomposition, but instead for features of a hidden state [18, 19]. Temporal memory (TM) algorithms can be used to make cumulative predictions about future states and their features, forming SR or SF. This work shows that the proposed algorithm, Distributed Hebbian Temporal Memory (DHTM), can dynamically form SFs for navigation tasks in Gridworld and AnimalAI [20] environments.
@@
The most prominent TM algorithms, such as a Recurrent Neural Network (RNN) [21] or LSTM [9], use backpropagation to capture data relationships, which is known to be unstable due to recurrent nonlinear derivatives. They also require complete data sequences to be available during training. Although the gradient vanishing problem can be partially circumvented in a way that Receptance Weighted Key Value (RWKV) [22] or Linear Recurrent Unit (LRU) [23] models do, the problem of online learning, i.e. learning on non-stationary data sets, is still a viable issue. Unlike graphical models such as HMM [24], RNN models and their descendants also lack a probabilistic theoretical foundation, which is advantageous for modelling sequences captured from stochastic environments [25, 26]. There is little research on TM models that can be used in fully online adaptive systems interacting with partially observable stochastic environments with access to only one sequence data point at a time, a common case in reinforcement learning [27].
@@
We propose a Distributed Hebbian Temporal Memory (DHTM) algorithm based on factor graph formalism and a multi-compartment neuron model inspired by the HTM neuron [28]. The resulting graphical structure of our model is similar to that of the factorial HMM [29]. An important feature of our model is that the transition matrices for each factor are stored as different components (segments) of artificial neurons, which makes computation very efficient in the case of sparse transition matrices. Our TM forms sequence representations entirely online and uses only local Hebbian-like learning rules [30, 31, 32], which avoid the drawbacks of gradient methods and make the learning process much faster than gradient methods. This comes at the expense of hidden state generalization, since DHTM’s hidden state uniquely encodes each observation sequence. Different trajectories leading to the same position in the partially observable Gridworld can result in different hidden states inferred by DHTM. In essence, this makes DHTM a model of episodic memory [33], or an adaptive trajectory buffer, with built-in machinery for making multistep future predictions. In this light, we also show that an agent forming SF with DHTM can be interpreted as a kind of episodic control [34, 35, 36, 37].
@@
The proposed TM is tested as a world model for an RL agent architecture navigating in a Gridworld environment and a more challenging AnimalAI testbed [38]. The results demonstrate that our algorithm outperforms classical LSTM and a biologically inspired HMM-like world model CSCG [39] in changing environments and is in order of magnitude more sample efficient than Dreamer V3 [40]. Another advantage of our algorithm is that it can be implemented on neuromorphic processors using only local learning rules.
@@ 2 Background
This section provides basic information about some of the concepts necessary to follow the paper.
@@ 2.1 Reinforcement Learning
This paper considers decision-making in a partially observable environment, usually formalized as a partially observable decision process [13]. A POMDP is defined as a tuple $\mathcal{M}=(S,A,P,R,O,D,\gamma)$, where $S$—state space, $A$—action space, $P(s,a,s`)=Pr(s`\ |\ s,a)$—transition function, $R(s)$–reward function, O—observation space, $D(a,s`,o)=Pr(o\ |\ a,s`)$—sensor model and $\gamma\in[1,0)$—discount factor, given a transition $s,a\to s`$, where $s\in S$, $a\in A$, $o\in O$. If $S,A,O$ are finite, $P,D$ can be viewed as real-valued matrices, otherwise, they are conditional density functions. Here we consider deterministic rewards that depend only on the current state, i.e. $R(s):S\to\mathbb{R}$.
@@ 2.2 Successor Representation
Successor representations are such representations of hidden states from which we can linearly infer the state value given the reward function [15]. We assume that the observation and state spaces are discrete.
@@
An extension of SR for large or continuous state spaces that we use is Successor Features [18], which is based on a similar idea of value function decomposition, but in a feature space instead (see details in Appendix A).
@@ 2.3 Hidden Markov Model
A partially observable Markov process can be approximated by a Hidden Markov Model (HMM) with hidden state space $H$ and observation space $O$. $O$ is the same as in $\mathcal{M}$, but $H$ is generally not equal to $S$. Variables $H_{t}$ represent an unobservable (hidden) approximate state of the environment that evolves over time, and observable variables $O_{t}$ represent observations that depend on the same time-step state $H_{t}$, and $h_{t},o_{t}$ are corresponding values of these random variables. For simplicity, we assume that actions are fully observable and that information about them is contained in $H_{t}$ variables. For the process of length $T$ with state values $h_{1:T}=(h_{1},\ldots,h_{T})$ and $o_{1:T}=(o_{1},\ldots,o_{T})$, the Markov property yields the following factorization of the generative model:
@@
In the case of a discrete hidden state variable, a time-independent stochastic transition matrix can be learned with the Baum-Welch algorithm [41], a variant of the expectation maximization (EM) algorithm. To compute the statistics for the expectation step, it uses the forward-backward algorithm, which is a special case of the sum-product algorithm [42].
Since the transition matrix grows quadratically with the size of the hidden state space, it may become infeasible to train the HMM in large state spaces. Since it is possible that not all hidden states are dependent on each other, the Factorial HMM can mitigate this problem by partitioning the hidden state spaces into smaller independent subspaces, providing an efficient way to handle distributed representations.
We base DHTM on ideas from the Baum-Welch algorithm and the Factorial HMM, while adapting them for online learning.
2.4 Episodic Control
Episodic Control (EC) is a general computational approach to storing and reusing an agent’s experience, inspired by the notion of episodic memory [33] from psychology.
[…]
In this paper, we refer to EC to better explain the principle of the proposed algorithm for solving RL tasks.
@@ 3.1 Factor Graph Model
Distributed Hebbian Temporal Memory is based on the sum-product belief propagation algorithm in a factor graph (see Figure 1A). Analogous to Factorial HMM [44], we divide the hidden space $H$ into subspaces $H^{k}$. There are four sets of random variables (RV) in the model: $H^{i}_{t-1}$—latent variables representing hidden states from the previous time step (context), $H^{k}_{t}$—latent variables for the current time step, and $\Phi^{k}_{t}$–feature variables. All random variables have categorical distributions. RV state values are denoted by corresponding lowercase letters: $h^{i}_{t-1}$, $h^{k}_{t}$, $\varphi^{k}_{t}$, $o^{lm}_{t}$.
Each variable $\Phi^{k}_{t}$ is considered independent and has a separate graphical model for increased computational efficiency. However, in practice, hidden variables of the same time step are statistically interdependent. We introduce their interdependence through a segment computation trick that goes beyond the standard sum-product algorithm (see equation 6).
@@
We will discuss only the upper block of the graph in Figure 1A, which is DHTM itself. The lower block—an encoder—is described in Appendix C. The only requirement for the encoder is that its output should be represented as states of categorical variables (features) for the current observation.
@@
where $\Omega_{k}=\{h^{i}_{t-1}:i\in\omega_{k}\}$, $\omega_{k}={i_{1},\dots,i_{n}}$—set of previous time step RV indexes included in $F_{c}^{k}$ factor, $(n+1)$—factor size.
@@ 3.2 Neural Implementation
For computational purposes, we translate the factor graph to the Hebbian learning neural network architecture (for a biological interpretation of the model, see Appendix B). As shown in Figure 1B, each RV can be viewed as a set of spiking neurons representing the states of the RV, i.e. $p(h^{k}_{t})=p(c^{j}_{t}=1)$, where $j$–is the index of a neuron corresponding to the state $h^{k}_{t}$. Cell activity is binary $c^{j}_{t}\in\{0,1\}$ (spike/no spike), and the probability can be interpreted as a spike rate. Factors $F^{k}_{c}$ and $M^{i}_{t-1}$ can be represented as vectors, where elements are factor values for all possible combinations of RV states included in the factor. Let’s denote the elements of the vectors as $f_{l}$ and $m_{u}$ respectively, where $l$ corresponds to a particular combination of $k$ state values $h^{k}_{t},h^{i_{1}}_{t-1},\dots,h^{i_{n_{l}}}_{t-1}$, and $u$ indexes all neurons representing states of RVs from previous time steps.
We can now rewrite equation 3 as the following:
$p(h_{t}^{k})\propto\sum_{l\in\mathrm{seg}(j)}L_{l}f^{k}_{l},$ (4)
@@
where $\mathbb{I}$—indicator function, $\mathrm{col}(\varphi^{k}_{t})$ is a set of hidden states connected to the feature state $\varphi^{k}_{t}$ forming a column. The shape of the emission factor is inspired by the presumed columnar structure of the neocortex and has been shown to induce a sparse transition matrix in HMM [39].
where $w_{ul}$—synapse efficacy or neuron specificity for segment, such that ${w_{ul}=p(s_{l}=1|c^{u}_{t-1}=1)}$, $w_{l}=\frac{1}{n_{l}}\sum_{u}w_{ul}$—average synapse efficacy for segment $l$, and $n_{l}$-number of cells in segment’s receptive field.
where ${E_{l}=\log f_{l}+\log L_{l}}$, $\mathrm{cells}[H^{k}_{t}]$—indexes of cells representing states for the variable $H^{k}_{t}$. Here we also approximate the logarithmic sum with the $\mathrm{max}$ operation, inspired by the neurophysiological model of segment aggregation by cells [45].

### v4

@@ Abstract
This paper presents a novel approach to address the challenge of online sequence learning for decision making under uncertainty in non-stationary, partially observable environments. The proposed algorithm, Distributed Hebbian Temporal Memory (DHTM), is based on the factor graph formalism and a multi-component neuron model. DHTM aims to capture sequential data relationships and make cumulative predictions about future observations, forming Successor Features (SFs). Inspired by neurophysiological models of the neocortex, the algorithm uses distributed representations, sparse transition matrices, and local Hebbian-like learning rules to overcome the instability and slow learning of traditional temporal memory algorithms such as RNN and HMM. Experimental results show that DHTM outperforms LSTM, RWKV and a biologically inspired HMM-like algorithm, CSCG, on non-stationary data sets. Our results suggest that DHTM is a promising approach to address the challenges of online sequence learning and planning in dynamic environments.
@@ 1 Introduction
Modelling sequential data is one of the essential tasks in artificial intelligence as it has many applications, including decision-making and learning world models (Moerland et al., 2023), natural language processing (Min et al., 2021), conversational AI (Dwivedi et al., 2023), time series analysis (Eraslan et al., 2019), and video and music generation (Ji et al., 2020). One of the classical approaches to modelling sequential data is to form a representation that stores and condenses the most relevant information about a sequence, and to find a general transformation rule of this information through the dimension of time (Lipton et al., 2015; Harshvardhan et al., 2020; Mathys et al., 2011). We refer to the class of algorithms that use this approach as temporal memory (TM) algorithms because they essentially model the cognitive ability of complex living organisms to remember the past experience and make future predictions based on that memory (Hochreiter & Schmidhuber, 1997; Friston et al., 2016; 2018; Parr & Friston, 2017).
Temporal memories differ in their ability to capture data dependencies and the extent to which they can generalise their predictions to unseen sequences. There are two extremes known in psychology as cognitive maps Tolman (1948), which extract commonalities or schemas generalising experience, and episodic memory Tulving et al. (1972), which stores fragments of experience separately. They both can be viewed as temporal memories, however, the former is slower and require a lot of data to built up a useful schema, according to contemporary computational models based on HMM (George et al., 2021) and Transformers (Dedieu et al., 2024), while the latter is instant and is usually viewed as a data buffer (Zhang et al., 2021).
@@
In this paper, we explore advantages of an episodic-like memory in addressing the problem of online sequence learning for planning and decision-making under uncertainty, which can be formalized as reinforcement learning (RL) for a partially observable Markov decision process (POMDP) (Poupart, 2005; Singh et al., 2021). Inferring the hidden state in a partially observable environment is a sequence modelling problem, since it requires processing a sequence of observations to obtain enough information about the hidden state. One of the most efficient representations of hidden states for discrete POMDPs is the Successor Representation (SR), which disentangles hidden states and goals given by the reward function (Dayan, 1993; Samuel J. Gershman, 2018; Millidge & Buckley, 2022). The Successor Features framework is an extension of SR to continuous POMDP, using the same idea of value function decomposition, but instead for features of a hidden state (Barreto et al., 2017; Touati et al., 2022). Temporal memory (TM) algorithms can be used to make cumulative predictions about future states and their features, forming SR or SF. This work shows that the proposed algorithm, Distributed Hebbian Temporal Memory (DHTM), can dynamically form SFs for navigation tasks in Gridworld and AnimalAI (Fountas et al., 2020) environments.
@@
The most prominent TM algorithms, such as a Recurrent Neural Network (RNN) (Boden, 2002) or LSTM (Hochreiter & Schmidhuber, 1997), use backpropagation to capture data relationships, which is known to be unstable due to recurrent nonlinear derivatives. They also require complete data sequences to be available during training. Although the gradient vanishing problem can be partially circumvented in a way that Receptance Weighted Key Value (RWKV) (Peng et al., 2023) or Linear Recurrent Unit (LRU) (Orvieto et al., 2023) models do, the problem of online learning, i.e. learning on non-stationary data sets, is still a viable issue. Unlike graphical models such as HMM (Eddy, 2004), RNN models and their descendants also lack a probabilistic theoretical foundation, which is advantageous for modelling sequences captured from stochastic environments (Salaün et al., 2019; Zhao et al., 2020). There is little research on TM models that can be used in fully online adaptive systems interacting with partially observable stochastic environments with access to only one sequence data point at a time, a common case in reinforcement learning (Jahromi et al., 2022).
@@
We propose a Distributed Hebbian Temporal Memory (DHTM) algorithm based on factor graph formalism and a multi-compartment neuron model inspired by the HTM neuron (Hawkins & Ahmad, 2016). Factor graph is a flexible graphical representation of probabilistic distributions, which is commonly used in combination with sum-product algorithm (Loeliger, 2004). We found this representation particularly convenient in our attempt of probabilistic interpretation of the HTM framework. The resulting graphical structure of our model is similar to that of the factorial HMM (Ghahramani & Jordan, 1995). An important feature of our model is that the transition matrices for each factor are stored as different components (segments) of artificial neurons, which makes computation very efficient in the case of sparse transition matrices. Our TM forms sequence representations entirely online and uses only local Hebbian-like learning rules (Hebb, 2005; Churchland & Sejnowski, 1992; Lillicrap et al., 2020), which avoid the drawbacks of gradient methods and make the learning process much more sample efficient than gradient methods. This comes at the expense of hidden state generalization, since DHTM’s hidden state uniquely encodes each observation sequence. Different trajectories leading to the same position in the partially observable Gridworld can result in different hidden states inferred by DHTM. In essence, this makes DHTM a model of episodic memory, or an adaptive trajectory buffer, with built-in machinery for making multistep future predictions. In this light, we also show that an agent forming SF with DHTM can be interpreted as a kind of episodic control (EC) (Lengyel & Dayan, 2007; Pritzel et al., 2017; Ritter et al., 2018; Emukpere et al., 2023).
Though our approach is similar to EC, we are not aware of other works, demonstrating its application to partially observable environments. There are other temporal memories that, similarly to DHTM, use Hebbian rules and learn online, like Asymmetric Hopfield Network (AHN) (Chaudhry et al., 2023), Temporal Predictive Coding (tPC) (Tang et al., 2023) and Hierarchical Temporal Memory (Hawkins & Ahmad, 2016). Compared to DHTM, AHN is a first-order memory. That is, it is not able to disentangle identical observations that correspond to different underlying environmental states. Therefore, it is not suitable for partially observable environments. In contrast, tPC has a hidden layer that, potentially, makes it a high-order memory. Despite that it has been tested only in MDP environments (Tang et al., 2024) so far, this method paves a promising direction, which we aim to compare with in future work. Finally, the algorithm that inspired our work, HTM, is also a Hebbian high-order temporal memory, which, opposed to DHTM, doesn’t allow for probabilistic predictions due to its all-or-nothing computations.
@@
The proposed TM is tested as an episodic memory for an RL agent architecture navigating in a Gridworld environment and a more challenging AnimalAI testbed (Crosby et al., 2020). The results demonstrate that our algorithm outperforms classical LSTM, RWKV and a biologically inspired HMM-like world model CSCG (George et al., 2021) in changing environments and is in order of magnitude more sample efficient than Dreamer V3 (Hafner et al., 2023). Another advantage of our algorithm is that it can be implemented on neuromorphic processors using only local learning rules.
@@ 2.1 Reinforcement Learning
This paper considers decision-making in a partially observable environment (Poupart, 2005). The environment is defined by its state transition function $P(s,a,s^{\prime})=Pr(s^{\prime}\ |\ s,a)$, determining its dynamics, where $s^{\prime},s\in S$ is called state space, $a\in A$—action space. Partial observability means that an agent has access to observations $o\in O$, which only partially inform about the environment’s state $s$ at each moment.
@@ 2.2 Successor Representation
Successor representations are such representations of hidden states from which we can linearly infer the state value given the reward function (Dayan, 1993). For a discrete state space, such representation can be derived just by rearranging sums in the equation for value function:
@@
An extension of SR for large or continuous state spaces that we use is Successor Features (Barreto et al., 2017), which is based on a similar idea of value function decomposition, but in a feature space instead (see details in Appendix A).
@@ 2.3 Hidden Markov Model
A partially observable environment can be approximated by a Hidden Markov Model (HMM) with hidden state space $H$ and observation space $O$. Generally, $H$ is not equal to environemnts state space $S$ since the environemnt’s true structure is usually unknown. For simplicity, we assume that actions are fully observable and that information about them is contained in hidden state. For the process of length $T$ with state values $h_{1:T}=(h_{1},\ldots,h_{T})$ and observations $o_{1:T}=(o_{1},\ldots,o_{T})$, the Markov property yields the following factorization of the generative model:
@@
In the case of a discrete hidden state variable, a time-independent stochastic transition matrix can be learned with the Baum-Welch algorithm (Baum et al., 1970), a variant of the expectation maximization (EM) algorithm. To compute the statistics for the expectation step, it uses the forward-backward algorithm, which is a special case of the sum-product algorithm (Kschischang et al., 2001).
Factor graph is a convenient graphical representation (Loeliger, 2004) of probabilistic models like HMM (see Fig.1A for an example). Formally, a factor graph is a bipartite graph, where one part consists of factors (squares), which, replacing conditional probabilities, define dependencies between random variables (circles) constituting the other part. Such representation also allows extension of forward-backward algorithm to sum-product algorithm. As shown in section 3.1, we base DHTM on ideas from the sum-product algorithm and the Factorial HMM, while adapting them for online learning.
@@ 3.1 Factor Graph Model
Distributed Hebbian Temporal Memory is based on the sum-product belief propagation algorithm in a factor graph (see Figure 1A). Analogous to Factorial HMM (Ghahramani & Jordan, 1997), we divide the hidden space $H$ into subspaces $H^{k}$. There are four sets of random variables (RV) in the model: $H^{i}_{t-1}$—latent variables representing hidden states from the previous time step (context), $H^{k}_{t}$—latent variables for the current time step, and $\Phi^{k}_{t}$–feature variables. All random variables have categorical distributions. RV state values are denoted by corresponding lowercase letters: $h^{i}_{t-1}$, $h^{k}_{t}$, $\varphi^{k}_{t}$, $o^{lm}_{t}$.
In this work, each variable $\Phi^{k}_{t}$ is considered independent and has a separate graphical model for increased computational efficiency. However, in practice, hidden variables of the same time step are statistically interdependent. We introduce their interdependence through a segment computation trick that goes beyond the standard sum-product algorithm (see Appendix B for details).
@@
We will discuss only the upper block of the graph in Figure 1A, which is DHTM itself. The lower block—an encoder—is described in Appendix D. The only requirement for the encoder is that its output should be represented as states of categorical variables (features) for the current observation.
@@
where summation is over all possible hidden RV states in $\Omega_{k}=\{h^{i}_{t-1}:i\in\omega_{k}\}$, $\omega_{k}={i_{1},\dots,i_{n}}$—set of previous time step RV indexes included in $F_{c}^{k}$ factor, $(n+1)$—factor size.
@@ 3.2 Neural Implementation
For computational purposes, we translate the factor graph shown in Figure 1A to the Hebbian learning neural network architecture shown in Figure 1B (for a biological interpretation of the model, see Appendix C). Each RV can be viewed as a set of neurons representing the states of the RV, i.e. $p(h^{k}_{t})=p(c^{j}_{t}=1)$, where $j$–is the index of a neuron corresponding to the state $h^{k}_{t}$ and $c^{j}_{t}\in\{0,1\}$—cell’s activity. Since RVs are categorical, only one cell per variable can be active at a time and their activity is defined through sampling of RV’s distribution.
Since factors $F^{k}_{c}$ represent dependencies between variables, in neural implementation we view them as collections of connections between cell groups corresponding to RVs that $F^{k}_{c}$ takes as arguments. Each factor’s entry maps a particular combination of states $h^{k}_{t},h^{i_{1}}_{t-1},\dots,h^{i_{n_{l}}}_{t-1}$ to a factor value $f_{l}$. We found it useful to represent each such entry as a separate computational unit, which we call segment (alluding to dendritic segments of a biological neuron). As shown in Figure 1B, segment $l$ connects its presynaptic cells $h^{i_{1}}_{t-1},\dots,h^{i_{n_{l}}}_{t-1}$ (one per each RV), constituting its receptive field, to the cell it excites $h^{k}_{t}$. The factor $f_{l}$ value can be thought of as a segment’s excitation efficiency, i.e. how strong its presynaptic pattern influences the cell’s activation.
Reformulation of belief propagation step in equation 3 in terms of segmental computations results in the following equations:
$p(h^{k}_{t})=\underset{j\in\mathrm{cells}[H^{k}_{t}]}{\mathrm{softmax}}(\underset{l\in seg(j)}{\mathrm{max}}(E_{l})),$ (4)
$E_{l}=\log f_{l}+w_{l}\log\frac{1}{n_{l}}\sum_{u\in\mathrm{rec}(l)}w_{ul}m_{u}+\sum_{u\in\mathrm{rec}(l)}(1-w_{ul})\log m_{u},$ (5)
where $\mathrm{cells}[H^{k}_{t}]$—indexes of cells corresponding to states of the variable $H^{k}_{t}$, $seg(j)$—all segments for cell $j$, $w_{ul}$—synapse efficacy or neuron’s specificity for segment, such that ${w_{ul}=p(s_{l}=1|c^{u}_{t-1}=1)}$, $w_{l}=\frac{1}{n_{l}}\sum_{u}w_{ul}$—average synapse efficacy for segment $l$, and $n_{l}$-number of cells in segment’s receptive field $rec(l)$. $m_{u}$ is presynaptic cell’s activation probability $p(c^{u}_{t-1}=1)$, which carries messages from the previous time-step and corresponds to entry in $M^{i}_{t-1}$. For detailed derivation and rationale behind these equations, see Appendix B.
Initially, cells have no segments. As learning progresses, new non-zero connections, grouped into segments, are created. In equation 5, we benefit from having sparse factors because their complexity depends linearly on the number of non-zero entries or segments. And that’s usually the case in our model due to the one-step Monte Carlo learning and the special form of the emission factors $F^{k}_{e}$:
@@
where $\mathbb{I}$—indicator function, $\mathrm{col}(\varphi^{k}_{t})$ is a set of hidden states connected to the feature state $\varphi^{k}_{t}$ forming a column. The shape of the emission factor is inspired by the presumed columnar structure of the neocortex and has been shown to induce a sparse transition matrix in HMM (George et al., 2021).
The next step after computing the $p(h^{k}_{t})$ distribution is to incorporate information about the current feature states ${p(h^{k}_{t}\ |\ \varphi^{k}_{t})\propto p(h^{k}_{t})\mathbb{I}[h^{k}_{t}\in\mathrm{col}(\varphi^{k}_{t})]}$. After that, the learning step is executed. The loop closing step of our TM algorithm is to assign the posterior of the current step ${p(h^{k}_{t}\ |\ \varphi^{k}_{t})}$ to $M^{i}_{t-1}$.